# StateFlow: Enhancing LLM Task-Solving through State-Driven Workflows

### v1

@@ Abstract
It is a notable trend to use Large Language Models (LLMs) to tackle complex tasks, e.g., tasks that require a sequence of actions and dynamic interaction with tools and environments. In this paper, we propose StateFlow, a novel LLM-based task-solving paradigm that conceptualizes complex task-solving processes backed by LLMs as state machines. With proper construction of states and definition of state transitions, StateFlow grounds the progress of task-solving, ensuring clear tracking and management of LLMs’ responses throughout the task-solving process. Within each state, StateFlow allows execution of a series of actions, involving not only the generation of LLM’s responses guided by a specific prompt, but also the utilization of external tools as needed. State transitions are controlled by specific rules or decisions made by the LLM, allowing for a dynamic and adaptive progression through the task’s pre-defined StateFlow model. Evaluations on the InterCode SQL and Bash benchmarks show that StateFlow significantly enhances LLMs’ efficiency.
@@ 1 Introduction
LLMs have increasingly been employed to solve complex, multi-step tasks, e.g., tasks that require a sequence of complex reasoning and interacting with external environments and tools Mialon et al. (2023); Chen et al. (2023b); Wang et al. (2023c); Yao et al. (2022). In approaching such complex tasks, there is typically a desired workflow, or plan of actions based on heuristics that could improve the efficiency of task solving Kim et al. (2023); Wu et al. (2023a). A common practice in the context of LLMs is to write a prompt to instruct the models to follow a specific procedure or to perform an action at a given condition. For example, Chain of thought, ReAct prompting, and the vast customization of GPTs Wei et al. (2022); Kojima et al. (2022); Yao et al. (2022). This relies on LLMs’ innate capability to determine the current task-solving status and perform subsequent actions autonomously. Despite the impressive abilities of LLMs, it is still unrealistic to expect LLMs to always make the correct judgment of the status of current progress. It is also almost impossible to reliably track these judgments and their decisions of subsequent action trajectory. Given these considerations, we pose the research question: How can we exert more precise control and guidance over large language models?
@@
In this paper, we propose StateFlow, a new framework which models LLM workflows as state machines. Finite State machines (FSMs) Mealy (1955); Moore et al. (1956) are used as control systems to monitor practical applications, such as traffic light control Wagner et al. (2006). A defined state machine is a model of behavior that decides what to do based on current status. A state represents one situation that the FSM might be in. Drawing from this concept, we want to use FSMs to model the task-solving process of LLMs. When using an LLM to solve a task with multiple reasoning steps, each step of the task-solving process can be mapped to a state. For example, the process starts at the Init state when the task is given. When reaching a state in StateFlow, a sequence of output functions is called to add content to the context history, including sending a specific instruction, using a tool, or calling the LLM itself. Based on current state and context history, the StateFlow model would determine the next state to transit to. The task-solving process progresses by transitioning through different states and perform corresponding actions, and ends until a final state is reached. Essentially, we are sending different instructions to an LLM to ask it to perform different actions based on current status. StateFlow enhances control over the task-solving process and can seamlessly integrate external tools and environments.
In StateFlow, we use FSM to control a single LLM, sending different instructions to it at different states. We also propose an agent view of our framework named SF_Agent, that can use different LLM agents to perform actions at different states. In this case, we don’t add the instructions to the context history. Instead, we construct individual agents with these instructions. This could reduce the length of the context history, and allow the flexibility to use different LLMs.
@@
Figure 1: The StateFlow models for the SQL and Bash task. The Init and End state are basic components of state machines, and states like Observe, Solve, Verify, Error are adaptable across various tasks. When reaching a state, a sequence of actions defined is executed. Execution outcomes are indicated by red arrows for failures and green for successes. At a success execution, there might be additional checks to determine the next state. For example, at a success ’Submit’ command, the model transits to End state.
We evaluate StateFlow on the SQL task and Bash task from the InterCode Yang et al. (2023a) benchmark. The results demonstrate the advantages of StateFlow over existing methods in terms of both effectiveness and efficiency. Additionally, we also further perform a case study on the SQL task to illustrate how to identify a desired workflow from a given task and build the StateFlow model. Based on the study, we modified the ReAct examples to follow the desired workflow, which resulted in an increase of 7% in success rate with GPT-3.5-Turbo. Building on the same workflow, our StateFlow model further outperforms the refined ReAct version by 6%, with 5$\times$ less cost.
@@
Our main contributions are the following: (1) We introduce StateFlow, a paradigm that models LLM workflows as state machines, allowing better control and efficiency in LLM-driven task solving. Building upon StateFlow, we introduce SF_Agent, an agent version of StateFlow, where LLM agents are used to perform specific actions when reaching a state. We also provide guidelines in how to build with the StateFlow framework and illustrate the building process through a case study. (2) We use two different tasks to illustrate the effectiveness and efficiency of StateFlow, with improvement in performance and a cost reduction of 4-6 times. We also perform additional ablations to provide deeper insights on how different states contribute to the performance of StateFlow.
@@ 2 Background
We first introduce state machines, which we will use to formulate our framework. A finite state machine (automaton) is a mathematical model of machine that accepts a set of words or string over an input alphabet $\Sigma$ Hopcroft et al. (2001); Carroll and Long (1989), where read of symbols would lead to state transitions. The automaton would determine whether the input is accepted or rejected. Formally, a Deterministic Finite-state Machine (DFSM) or a Deterministic Finite-state acceptor is a quintuple $\langle\Sigma,\mathrm{S},\mathrm{s}_{0},\delta,\mathrm{F}\rangle$ Rich et al. (2008), where:
  • $\Sigma$ is a finite non-empty set of symbols.
@@
  • F is the set of final states, a subset of $S$.
An acceptor is one of the simplest form of state machines, usually used as a language recognizer which determines whether an input ends in an accept state. This is not suitable in our modeling of LLM generations, where we want to determine actions to be performed and produce outputs in between states. Instead, we would base our model on a transducer finite-state machine, which is a sextuple $\langle\Sigma,\Gamma,\mathrm{S},\mathrm{s}_{0},\delta,\omega\rangle$, with $\mathrm{F}$ removed and two additional terms added:
  • $\Gamma$: is the output alphabet (a finite non-empty set of symbols).
  • $\omega$: is the output function.
Language Model Prompting
For an input-output prompting, given a probabilistic language model $p_{\theta}$, a question or task $Q$, and prompt $T$, our objective is to maximize the likelihood of obtaining correct answer A: $p(A\mid T,Q)=\prod_{i=1}^{|A|}p_{\theta}\left(a_{i}\mid T,Q,a_{<i}\right)$, where the $i$-th token is predicted based on the prompt, question, and previous tokens Qiao et al. (2022). Chain of Thoughts (CoT) Wei et al. (2022) generates a sequence of intermediate steps, and the final answer is derived based on all previous steps. We denote $C$ as a sequence of intermediate reasoning steps $C_{1},C_{2},...C_{j}$. We can formalize CoT prompting to be: $p(A\mid T,Q)=p(A\mid T,Q,C)p(C\mid T,Q)$, where $p(C\mid T,Q)=\prod_{j=1}^{|C|}p_{\theta}\left(C_{j}\mid T,Q,C_{<j}\right)$, represents the generation of sequential intermediate steps. In terms of tool use, each intermediate step may have an initial thought $C_{i}$ to utilize a tool, and the feedback $O_{i}$ is concatenated back.
@@ 3 Methodology
In this section, we first define the StateFlow model and an agent version SF_Agent. We then provide a general guideline for constructing StateFlow model for any task with a detailed case study.
@@ 3.1 StateFlow
In a finite state machine, a state is a representation of its past history, or an implicit representation of how the state machine has reached the present situation Wagner et al. (2006). We think it is feasible to conceptualize the process of LLM inference as a state machine, using states to represent the context history. While a transducer state machine transits between states and generate outputs based on current state and an input tape, StateFlow doesn’t have the concept of input tape but solely depends on the context history, which is a cumulative record of all past interactions. StateFlow employs a set of instructions $T={T_{1},T_{2},..T_{j}}$ to guide the language model generation through different states. A prompt $T_{i}$ is issued only when a condition is met (reaching a given state). This dynamic prompting approach ensures that the language model receives the most relevant guidance at each state, enhancing its ability to tackle complex problems.
Following the definition of finite state machines, we formulate a StateFlow model to be a sextuple $\langle\mathrm{S},\mathrm{s}_{0},\mathrm{F},\delta,\Gamma,\Omega\rangle$ and explain each of the components under the LLM scenario:
States $\mathrm{S}$
@@
A state represents the current status of a running process, essentially abstracting the historical context. Upon entering a state, a predefined set of actions is executed. For example, entering an error state implies the process encounters an issue, prompting the execution of predetermined error-handling actions.
Initial state $\mathrm{s}_{0}$
@@
The process begins at the initial state when receiving the input task/question.
Final States $\mathrm{F}$
@@
A set of final states when the process is terminated, which is a subset of $\mathrm{S}$.
Output $\Gamma$
@@
We define $\Gamma$ to be an infinite set of messages (unit of text) consisting of prompts $T$, language model responses $C$, and tool calls $O$: $\Gamma=\{T,C,O\}$, which represents all possible messages that can be generated within StateFlow. We further define context history to be $\Gamma^{*}$, which is a list of messages that have been generated. $S\times\Gamma^{*}$ could be view as a snapshot of a running StateFlow Rich et al. (2008).
State transition $\delta$
@@
Based on the current state and context history, the transition function would determine which state to go to ($\delta:S\times\Gamma^{*}\rightarrow S$). This could mean string matching of the last input, for example, checking if ‘Error’ is in the message from code execution. We can also explicitly employ an LLM to check for conditions and determine what is the next state.
Output Functions $\Omega$
Here we define $\Omega$ to be a set of output functions, where a function takes context history and generates an output ($\Omega:\Gamma^{*}\rightarrow\Gamma$). It is either an LLM, a tool call, or a prompter function that returns a predefined prompt. The generated response will then be added to the context history.
Differing from traditional FSMs, we don’t have a set of symbols as input. We define a set of output functions $\Omega$ (instead of a single $\omega$) to distinguish different types of outputs. These output functions all depend on the context history, for example, LLMs perform next-token prediction based on previous context, and a tool reads the past conversation (usually the last message) to generate feedback. The process starts at $s_{0}$ when task $Q$ is appended to the context history $\Gamma^{*}$ and ends when reaching one of the final states (See Algorithm 1). We also use a counter that defines maximum turns of transitions to prevent infinite loops. The process returns the exit state and the whole context history in the end.
SF_Agent
Through constructing individual LLM agents with a specific prompt, we identify SF_Agent as a variation of StateFlow, where the workflow is built through defining the sequence of calling to different agents and tools. In this case, $\Gamma=\{C,O\}$, without instructions $T$. We would also remove the prompter from the output functions, but has a set of specialized agents $p_{\theta}^{T_{i}}$, where $T_{i}$ is a single instruction.
@@
3.2 StateFlow Design
We now provide a guideline for defining a StateFlow model (we refer to Section 3.3 for a case study). In general, StateFlow is more useful for tasks that require a designated process to solve. Designing a state machine is translating conceptual control flow or a human’s reasoning process into a logical model, based on one’s understanding of a given task.
Defining States
To define the states, we start with identifying an ideal workflow of a given task, or the Sunny-Day-Scenario that everything goes well. A state should represent a distinct phase or step in the process, defined with enough granularity to capture key milestones and decision points. With the basic workflow in mind, we need to think about possible situations during the process. Specifically, handling failures is an important part of the state machine, where a Error state is commonly used to handle failures. A set of fine-grained states might lead to better control over a problem-solving process. For example, it is possible to identify different types of errors and use different ways to handle them, but this adds complexity in defining the model as a trade-off.
Defining output functions
Within each state, we need to define a set of outputs. In practice, we need to identify the set of tools we will use, and what instructions we should send to the LLM. For example, in the Solve state defined for the bash task, we first send an instruction that asks for a bash command, call the model, and then execute the bash command. This is a typical sequence of sending instructions to LLMs and then utilizing tools.
Defining Transitions
We identify two possible state transitions: 1. We can use a static string matching with the LLM responses or tool executions. For example, the tool execution might return a specific string such as "execution failed", which can be used to determine the state transition. Also, LLMs are usually instructed to follow a certain template to generate outputs, so it is also possible to extract strings from the LLM response for the state transition. 2. When the previous context is too random to be used for string matching, we can always call LLM for an explicit checking of the current status. For example, we can send the context history to a model to ask whether the problem is solved.
@@ 3.3 A Case Study of StateFlow Design on SQL
We use the SQL task from the InterCode Yang et al. (2023a) benchmark to illustrate how to identify a workflow and convert it to a StateFlow model. The benchmark is evaluated with ReAct Yao et al. (2022), a few-shot prompting that requires annotations of example trajectories of actions, thoughts, and observations, which demonstrate the desired pattern to interact with tools or external feedback. These trajectories are human-crafted, and they reflect how one thinks is a good process to solve the problem. There are 4 trajectories in the prompt, and please see Figure 2 for one trajectory. It demonstrates what to do based on previous history, which is similar to state transitions in StateFlow: 1. the process starts with a ‘SELECT’ query and results in an error. 2. At an error, the next thought is to explore the tables, so the SHOW TABLES command is executed to retrieve all tables. 3. After getting the tables, the next step is to explore the ‘highschoolers’ table with the ‘DESC’ command. 4. With knowing what the table contains, the next thought is to use the select query to solve the question. 5. Finally, it confirms the output contains relevant info, and submit. We can see that the next thought is based on the status of the current solving process. While the 4 trajectories all start with a ‘SELECT’ query but result in an error, we believe a better workflow would be to first explore the tables and use the ‘SELECT’ command. Based on this, we construct the states to be: Init -> Observe -> Solve -> Verify -> End (See Figure 1). In each state, we instruct the model to perform a specific action, for example, we ask the model to submit if the task is verified to be solved in state Verify. We also have an error state, that we instruct the model to explore tables at an error.

### v2

@@ Abstract
It is a notable trend to use Large Language Models (LLMs) to tackle complex tasks, e.g., tasks that require a sequence of actions and dynamic interaction with tools and external environments. In this paper, we propose StateFlow, a novel LLM-based task-solving paradigm that conceptualizes complex task-solving processes as state machines. In StateFlow, we distinguish between “process grounding” (via state and state transitions) and “sub-task solving” (through actions within a state), enhancing control and interpretability of the task-solving procedure. A state represents the status of a running process. The transitions between states are controlled by heuristic rules or decisions made by the LLM, allowing for a dynamic and adaptive progression. Upon entering a state, a series of actions is executed, involving not only calling LLMs guided by different prompts, but also the utilization of external tools as needed. Our results show that StateFlow significantly enhances LLMs’ efficiency. For instance, StateFlow achieves 13% and 28% higher success rates compared to ReAct in InterCode SQL and ALFWorld benchmark, with 5$\times$ and 3$\times$ less cost respectively. We also show that StateFlow can be combined with iterative refining methods like Reflexion to further improve performance.
@@ 1 Introduction
LLMs have increasingly been employed to solve complex, multi-step tasks. Specifically, they have been applied to tasks that require interactions with environments (Yang et al., 2023a; Yao et al., 2022a; Shridhar et al., 2020) and those tasks that can benefit from utilizing tools such as web search and code execution (Mialon et al., 2023; Wu et al., 2023b; Wang et al., 2023c). In approaching these tasks, there is typically a desired workflow, or plan of actions based on heuristics that could improve the efficiency of task solving (Kim et al., 2023; Wu et al., 2023a). A common practice in the context of LLMs, such as ReAct (Yao et al., 2022b) and the vast customization of GPTs, is to write a single prompt that instructs the models to follow a desired procedure to solve the task. The LLM is called iteratively with the same instruction, along with previous actions and feedback from tools/environments. This relies on LLMs’ innate capability to determine the current task-solving status and perform subsequent actions autonomously. Despite the impressive abilities of LLMs, it is still unrealistic to expect LLMs to always make the correct judgment of the status of current progress. It is also almost impossible to reliably track these judgments and their decisions of subsequent action trajectory. Given these considerations, we pose the research question: How can we exert more precise control and guidance over LLMs?
@@
In this paper, we propose StateFlow, a new framework that models LLM workflows as state machines. Finite State Machines (FSMs) (Mealy, 1955; Moore et al., 1956) are used as control systems to monitor practical applications, such as traffic light control (Wagner et al., 2006). A defined state machine is a model of behavior that decides what to do based on current status. A state represents one situation that the FSM might be in. Drawing from this concept, we want to use FSMs to model the task-solving process of LLMs. When using LLMs to solve a task with multiple steps, each step of the task-solving process can be mapped to a state. For example, to solve the InterCode (Yang et al., 2023a) SQL task, a desired procedure is to first gather information about the tables and columns in the database, then construct a query to retrieve required information, and finally verify the task is solved and end the process. We can convert this workflow to a set of states (See upper left of Figure 1). Within each state, we define a sequence of output functions, which will be called upon entering the state. The output functions would take in the context history and output a new context to be appended to the history, which can be a tool, an LLM with a specific instruction, or a prompter. Based on the current state and context history, the StateFlow model would determine the next state to transit to. The task-solving process progresses by transitioning through different states and calls to corresponding output functions, and ends until a final state is reached. Thus, StateFlow enhances control over the task-solving process and seamlessly integrates external tools and environments.
@@
Figure 1: The StateFlow models for the SQL and Bash task. Init and End state are basic components of state machines, and states like Observe, Solve, Verify, Error can be adaptable across various tasks. When reaching a state, a sequence of output functions defined is executed (e.g., $\text{M}_{i}\rightarrow\text{E}$ means to first call the model and then call the SQL/Bash execution). Execution outcomes are indicated by red arrows for failures and green for successes. Transition to different states is based on specific rules. For example, at a success ‘Submit’ command, the model transits to End state.
We evaluate StateFlow on the SQL task and Bash task from the InterCode (Yang et al., 2023a) benchmark and the ALFWorld Benchmark (Shridhar et al., 2020). The results demonstrate the advantages of StateFlow over existing methods in terms of both effectiveness and efficiency. With GPT-3.5, StateFlow outperforms ReAct by 13% on InterCode SQL task and 28% on ALFWorld, with 5$\times$ and 3$\times$ less LLM inference cost respectively. StateFlow is orthogonal to methods that iteratively improve future attempts using feedback based on previous trials (Shinn et al., 2023; Madaan et al., 2024; Prasad et al., 2023). Notably, we show that StateFlow can be combined with Reflexion (Shinn et al., 2023), improving the success rate on ALFWorld from 84.3% to 94.8% after 6 iterations.
@@
Our main contributions are the following: (1) We introduce StateFlow, a paradigm that models LLM workflows as state machines, allowing better control and efficiency in LLM-driven task solving. We provide guidelines on how to build with the StateFlow framework and illustrate the building process through a case study. (2) We use three different tasks to illustrate the effectiveness and efficiency of StateFlow, with improvement in performance and a 3-5$\times$ cost reduction. We also perform an ablation study to provide deeper insights into how different states contribute to the performance of StateFlow. (3) We show that StateFlow can be combined with iterative refining methods to further improve performance.
@@ 2 Background
We first introduce state machines, which we will use to formulate our framework. A finite state machine (automaton) is a mathematical model of a machine that accepts a set of words or string over an input alphabet $\Sigma$ (Hopcroft et al., 2001; Carroll & Long, 1989), where read of symbols would lead to state transitions. The automaton would determine whether the input is accepted or rejected. A basic Deterministic Finite-state Machine (DFSM) is a Deterministic Finite-state acceptor, usually used as a language recognizer that determines whether an input ends in an accept state. An acceptor is not suitable in our modeling of LLM generations, where we want to determine actions to be performed and produce outputs in between states. Instead, we base our model on a transducer finite-state machine, which is a sextuple $\langle\Sigma,\Gamma,\mathrm{S},\mathrm{s}_{0},\delta,\omega\rangle$ (Rich et al., 2008):
  • $\Sigma$ is the input alphabet (a finite non-empty set of symbols).
  • $\Gamma$: is the output alphabet (a finite non-empty set of symbols).
@@
  • F is the set of final states, a (possibly empty) subset of $S$.
@@ 3 Methodology
In this section, we first define the StateFlow model. We then provide a general guideline for constructing StateFlow model and illustrate with a detailed case study.
@@ 3.1 StateFlow
In a finite state machine, a state carries information about the machine’s history, tracking how the state machine has reached the present situation (Wagner et al., 2006). We think it is feasible to conceptualize the task-solving process with LLMs as a state machine, and define states and transitions to model the various stages throughout the process. While a transducer state machine transits between states and generates outputs based on the current state and an input tape, StateFlow doesn’t have the concept of input tape but solely depends on the context history, which is a cumulative record of all past interactions. StateFlow employs a set of instructions $T={T_{1},T_{2},..T_{i}}$ to guide the language model generation at different states. This is equivalent to constructing a set of LLM agents $p_{\theta}^{T_{i}}$ with a specific instruction $T_{i}$. This dynamic prompting approach ensures that the language model receives the most relevant guidance at each state, improving its ability to focus on a specific step. Following the definition of finite state machines, we formulate a StateFlow model to be a sextuple $\langle\mathrm{S},\mathrm{s}_{0},\mathrm{F},\delta,\Gamma,\Omega\rangle$ and explain each of the components under the LLM scenario:
@@
States $\mathrm{S}$   A state encapsulates the current status of a running process, essentially an abstraction of the context history. Upon entering a state, a predefined set of actions is executed. For example, entering an error state implies the process encounters an issue, triggering the execution of predetermined error-handling actions.
@@
Initial state $\mathrm{s}_{0}$  The process begins at the initial state when receiving the input task/question.
@@
Final States $\mathrm{F}$  A set of final states when the process is terminated, which is a subset of $\mathrm{S}$.
@@
Output $\Gamma$  We define $\Gamma$ to be an infinite set of messages (unit of text) consisting of prompts $P$, language model responses $C$, and feedback from tool/environment $O$: $\Gamma=\{P,C,O\}$, which represents all possible messages that can be generated within StateFlow. We further define context history to be $\Gamma^{*}$, which is a list of messages that have been generated. $S\times\Gamma^{*}$ could be view as a snapshot of a running StateFlow (Rich et al., 2008). Here we distinguish static prompts $P$ that will added to the context directly from instructions $T$ that are used upon calling an LM.
@@
State transition $\delta$  Based on the current state and context history, the transition function would determine which state to go to ($\delta:S\times\Gamma^{*}\rightarrow S$). This could mean string matching of the last input, for example, checking if ‘Error’ is in the message from code execution. We can also explicitly employ an LLM to check for conditions and determine what is the next state.
@@
Output Functions $\Omega$  Here we define $\Omega$ to be a set of output functions, where a function $\omega$ takes the whole context history and generates an output ($\omega:\Gamma^{*}\rightarrow\Gamma$). The output function can be an LLM, a tool call, or a prompter function that returns a static prompt (e.g., P, E and $\text{M}_{i}$ in Figure 1). The generated response will then be added to the context history.
Differing from traditional FSMs, we don’t have a set of symbols as input. We define a set of output functions $\Omega$ (instead of a single $\omega$) to distinguish different types of outputs. These output functions all depend on the context history, for example, LLMs perform next-token prediction based on the previous context, and a tool reads the past conversation (usually the last message) to generate feedback. The process starts at $s_{0}$ when task $Q$ is appended to the context history $\Gamma^{*}$ and ends when reaching one of the final states (See Algorithm 1). We also use a counter that defines maximum turns of transitions to prevent infinite loops. The process returns the exit state and the whole context history in the end.
3.2 Deployment Guideline
In this section, we provide guidelines for deploying StateFlow models, with reference to the case study discussed in Section 3.3. In general, StateFlow is designed for tasks that require a designated process to solve. Essentially, creating a state machine involves transforming abstract control flow or human reasoning into a formalized logical model, grounded in a comprehensive understanding of the task at hand.
Defining States   To define the states, we start with identifying an ideal workflow of a given task, or the Sunny-Day-Scenario that everything goes well. A state should represent a distinct phase or step in the process, defined with enough granularity to capture key milestones and decision points. With the basic workflow in mind, we need to think about possible situations during the process. Specifically, handling failures is an important part of the state machine, where a Error state is commonly used to handle failures. A set of fine-grained states might lead to better control over a problem-solving process. For example, it is possible to identify different types of errors and use different ways to handle them, but this adds complexity in defining the model as a trade-off.
Defining output functions  Within each state, we need to define a set of outputs. In practice, we need to identify the set of tools we will use, and what instructions we should send to the LLM. For example, in the Solve state defined for the bash task, we first send an instruction that asks for a bash command, call the model, and then execute the bash command. This is a typical sequence of sending instructions to LLMs and then utilizing tools.
Defining Transitions  We identify two possible state transitions: 1. We can use a static string matching with the LLM responses or tool executions. For example, the tool execution might return a specific string such as ”execution failed”, which can be used to determine the state transition. Also, LLMs are usually instructed to follow a certain template to generate outputs, so it is also possible to extract strings from the LLM response for the state transition. 2. When the previous context is too random to be used for string matching, we can always call LLM for an explicit checking of the current status. For example, we can send the context history to a model to ask whether the problem is solved.
@@ 3.3 A Case Study of StateFlow Design on SQL
We use the SQL task from the InterCode (Yang et al., 2023a) benchmark to illustrate how to identify a workflow and convert it to a StateFlow model. The benchmark is evaluated with ReAct (Yao et al., 2022b), a few-shot prompting that requires annotations of example trajectories of actions, thoughts, and observations. These trajectories are human-crafted, and they reflect how one thinks is a good process to solve the problem. See Figure 2 for one ReAct trajectory for the SQL task: (1) the process starts with a ‘SELECT’ query and results in an error. (2) At an error, the next thought is to explore the tables, so the SHOW TABLES command is executed to retrieve all tables. (3) After getting the tables, the next step is to explore the ‘highschoolers’ table with the ‘DESC’ command. (4) With knowing what the table contains, the next thought is to use the select query to solve the question. (5) Finally, it confirms the output contains relevant info, and submits. The trajectory demonstrates what to do based on previous history, which is similar to state transitions in StateFlow. While the trajectory starts with a ‘SELECT’ query but results in an error, we believe a better workflow would be to first explore the tables and use the ‘SELECT’ command. Based on this, we construct the states to be: Init -> Observe -> Solve -> Verify -> End (See Figure 1). In each state, the model is instructed to perform a specific action. For example, we ask the model to submit if the task is verified to be solved in state Verify. We also have an error state, that we instruct the model to explore tables at an error.
4 Experiments
4.1 InterCode Benchmark
We first experiment with 2 tasks from the InterCode benchmark (Yang et al., 2023a): (1) SQL: The InterCode-SQL adapts the Spider dataset for MySQL, containing 1034 task instances. For each task, a MySQL interpreter is set up with all relevant tables within a docker container. (2) Bash: The InterCode-Bash dataset has 200 task instances curated from the NL2Bash dataset. For both tasks, we allow a max of 10 rounds of interaction with the environment. We evaluate with OpenAI GPT-3.5-Turbo and GPT-4-Turbo (both with the 1106 version) and the temperature is set to 0.
Baselines. We compare StateFlow with two prompting strategies used in the InterCode benchmark. (1) Plan & Solve (Wang et al., 2023b): A two-step prompting strategy to first ask the model to propose a plan and then execute it. (2) ReAct (Yao et al., 2022b): a few-shot prompting method that prompts the model to generate thoughts and actions. Additionally, since we observed an ideal workflow for the SQL task in Section 3.3, we edit the ReAct prompt used in the benchmark accordingly to see if it already improves performance, named ReAct_Refined, and evaluate it on the SQL task (See Table 18 and 19 for details).
Metrics. We present metrics provided by the benchmark and we also report the LLM usage of each method. Success Rate (SR): a task is considered a success only if the reward is 1. Error Rate: percentage of commands that fails. Turns: number of interactions with the environment. Cost: the cost of calling an LLM API in US dollars.